The population of a country increased by an average of 2% per year from 2000 to 2003. If the population of this country was 2 000 000 on December 31, 2003, then the population of this country on January 1, 2000, to the nearest thousand would have been 

Grade:12

1 Answers

Harsh Patodia IIT Roorkee
8 years ago

Population in succeed year is 1.02 time previous year ( 2% is 2/100 i.e 1+0.02)
So population in 2003 will be 1.02(Population in 2002) = 1.02 ( 1.02(Poupluation in 2001)) =1.02*1.02*1.02(Population in 2001)
So population in 2001 is poulation in 2003/ (1.02) ​3
= 2000000/(1.02)​3